# Uni3C: Unifying Precisely 3D-Enhanced Camera and Human Motion Controls for Video Generation

### v1

@@ Discussion.
We empirically find that our method demonstrates robust performance even when working with imperfect point clouds obtained through monocular depth unprojection. In this context, the point clouds serve as the primary camera control signal, facilitating the convergence of training rather than dominating the multi-view geometric and textural generations, as illustrated in Figure 3. By achieving earlier model convergence, we can better maintain the inherent capabilities of VDMs.
4.2 Human Animation
In this paper, we explore the unified control through two human animation approaches, both of which are built on the Wan2.1 framework, targeting I2V and T2V, respectively. While these methods are not the primary focus of our work, we provide a brief introduction here. Formally, the concurrently pioneering work, RealisDance-DiT [73], is used to replace the Wan-I2V backbone for high-quality human animation. RealisDance-DiT incorporates SMPL-X [36] and Hamer [37] as additional input conditions to guide human motions. To ensure the flexibility for motion transfer, RealisDance-DiT randomly selects the reference frame in the video sequence, which is not perfectly aligned with the given SMPL-X. RealisDance-DiT only trains self-attention modules and patchify encoders to confirm the generalization. However, we clarify that integrating the control branch trained within different backbones is still challenging, which is addressed by our elaborate PCDController. Furthermore, we tried another version, RealisDance-DiT-T2V, based on Wan-T2V without reference image conditions to explore the generalization of PCDController. Remarkably, PCDController adapts successfully to Wan-T2V, empowering it with impressive I2V ability.
4.3 Global 3D World Guidance
Figure 4: The assignment of multi-modal conditions for Uni3C. Camera, point clouds, SMPL-X [36] and Hamer [37] are all presented as 3D conditions.
Figure 5: The overview of global 3D world guidance. (a) We first align the SMPL-X characters from the human world space $W_{hum}$ to the environment world space $W_{env}$ comprising dense point clouds. (b) GeoCalib [53] is used to calibrate the gravity direction of SMPL-X. (c) The rigid transformation coefficients $\tilde{s},\tilde{R},\tilde{t}$ are employed to align the whole SMPL-X sequence. We re-render all aligned conditions under specific camera trajectories as the global 3D world guidance.
@@ 4.3 Global 3D World Guidance
As illustrated in Figure 4, Uni3C adopts multi-modal conditions, including camera, point clouds, reference image, SMPL-X [36], and Hamer [37]. The first three conditions are used for PCDController, while the latter three are used for human animations. We employ GVHMR [48] to recover SMPL-X characters. One can also retrieve desired motion sequences from motion datasets [39, 41, 16, 31] or generate new motions through text-to-motion models [24, 15]. Although most of the conditions above are formulated as 3D presentations, they stay in two different world coordinates. We define the point cloud world coordinate as the environmental world space $W_{env}$, which is the main world space controlled by cameras, while the SMPL-X is placed in the human world space $W_{hum}$. It is non-trivial to control the camera across two different world spaces consistently. For example, determining the initial camera position within $W_{hum}$ is particularly ambiguous, especially for tasks involving motion transfer. Therefore, we propose the global 3D world guidance that places the human condition into the “environment”, i.e., aligning SMPL-X from $W_{hum}$ to $W_{env}$ as shown in Figure 5(b).
@@
Fortunately, 2D human pose keypoints subtly bridge $W_{hum}$ and $W_{env}$. Formally, we first estimate 17 human keypoints $\{\mathbf{k}^{i}_{2D}\}_{i=1}^{17}\in\mathbb{R}^{17\times 2}$ from the reference view $I_{img}$ by ViTPose++ [61]. Then, we unproject 2D keypoints into $W_{env}$ to obtain 3D keypoints $\{\mathbf{k}^{i}_{env}\}_{i=1}^{17}\in\mathbb{R}^{17\times 3}$ through metric monocular depth $\hat{D}_{img}$ and the intrinsic camera of the reference image. For SMPL-X in $W_{hum}$, the COCO17 regressor [48] is utilized to project the first SMPL-X character into $\{\mathbf{k}^{i}_{hum}\}_{i=1}^{17}$ corresponding to the same human keypoints. Consequently, a least-squares estimation [52] based rigid transformation can be used to align $\mathbf{k}_{hum}$ to $\mathbf{k}_{env}$ as follows:
$\min_{\tilde{s},\tilde{R},\tilde{t}}\sum_{i=1}^{17}w_{i}\|(\tilde{s}\tilde{R}(\mathbf{k}^{i}_{hum})^{T}+\tilde{t})^{T}-\mathbf{k}^{i}_{env}\|^{2},$ (4)
@@
where $\tilde{s},\tilde{R},\tilde{t}$ indicate the optimized scaling factor, rotation matrix, and translation vector, respectively. $w_{i}$ denotes the confidence weight of 2D keypoint $\mathbf{k}^{i}_{2D}$. We discard any keypoints with confidence below 0.7, as they typically degrade alignment quality. Once the transformation parameters $\tilde{s},\tilde{R},\tilde{t}$ are determined, we apply them to all other SMPL-X sequences under the assumption that they share the same rigid transformation. However, even minor orientation errors can accumulate, leading to physically unrealistic motion trajectories, such as ascending into the sky or descending into the ground. To address this, we adopt GeoCalib [53] to estimate the gravity direction in $W_{env}$, which is then employed to calibrate the SMPL-X to ensure parallel gravity directions, as illustrated in Figure 5(b). For the alignment of Hamer [37], which shares common vertices with the hand parts of SMPL-X, Hamer can also be aligned to $W_{env}$ through the rigid transformation (Equation 4). Additionally, we address the issue of hand occlusion for Hamer by masking occluded hand parts based on the rendered depth from SMPL-X. After the alignments of SMPL-X and Hamer sequences, we place all conditions into $W_{env}$, establishing the global 3D world guidance that allows for rendering concurrently controlled conditions under arbitrary camera trajectories and human motions as shown in Figure 5(c). Finally, these re-rendered conditions are sent to PCDController and RealisDance-DiT for generated outcomes, as shown in Figure 4.
5 Experiments
5.1 Implementation Details
@@
PCDController is trained with the frozen Wan-I2V [54] on multi-resolution images scaled from [$480\times 768$, $512\times 720$, $608\times 608$, $720\times 512$, $768\times 480$] of 81 frames. The learning rate is warmed up to 1e-5 for 400 steps and then fixed. We train the model for 6,000 steps with a batch size of 32, while more training steps would slightly hurt Wan-I2V’s generalization. The training is accomplished with 64 H100 GPUs for 40 hours. We also provided results based on CogVideoX-5B-I2V [64], training for 20k steps with a batch size of 16. During the training, we randomly drop 10% texts, as well as 5% point cloud renderings and Plücker embeddings. For inference, we set the classifier-free guidance scale to 5.0 for textual conditions, keeping other guidance on the default scale 1.0.
@@
To ensure the generalization of PCDController, we collect large-scale training data for camera control, including DL3DV [32], RE10K [74], ACID [34], Co3Dv2 [43], Tartainair [56], Map-Free-Reloc [1], WildRGBD [59], COP3D [49], UCo3D [35]. This comprehensive dataset encompasses a variety of scenarios, featuring both static and dynamic scenes, as well as object-level and scene-level environments. Furthermore, all datasets are annotated with metric-aligned monocular depth through the way proposed in [9] or are provided with ground-truth depth.
5.2 Results of Camera Control
Table 1: Quantitative results of camera control. VBench++ scores (%) are normalized (higher is better). Injected camera features are divided as Plücker ray and point clouds (Pcd). $\dagger$ denotes the results with challenging 360 camera rotations.
@@ 5.2 Results of Camera Control
ViewCrafter [66] $\checkmark$ 85.39 89.69 91.68 55.13 64.33 92.94 97.66 95.59 96.11 0.210 0.117 0.873
SEVA [71] $\checkmark$ 87.39 91.86 93.37 56.79 68.43 95.74 98.59 97.08 97.23 0.077 0.029 0.223
@@
PCDController $\checkmark$ $\checkmark$ 88.27 92.20 93.37 58.99 71.96 95.56 98.66 97.38 98.01 0.102 0.031 0.246
Ours (Wan-I2V)$\dagger$ $\checkmark$ 1.327 0.551 6.334
@@
To evaluate camera control ability, we build an out-of-distribution benchmark with 32 images across various domains, including text-to-image generation [^1] https://github.com/black-forest-labs/flux, real-world [25, 4], object-centric, and human scenes as displayed in Figure 8. Each image features four distinct camera trajectories, resulting in 128 test samples. We used VBench++ [23] to evaluate the video quality, while absolute translation error (ATE), relative translation error (RPE), and relative rotation error (RRE) are used to verify the camera precision. We use VGGT [55] to produce the extrinsic cameras of generated images, which are evaluated with the pre-defined cameras after the trajectory alignment.
@@
We present the quantitative results in Table 1, comparing our model with ViewCrafter [66], SEVA [71], and the CogVideoX [64] version of our framework. The qualitative outcomes are shown in Figure 9. Our experiments demonstrate that point clouds significantly enhance the controllability of both the Wan2.1 and CogVideoX, as verified by the improvement of ATE, RPE, and RRE. Although SEVA achieves precise camera trajectories, it requires massive training with static multi-view data (0.8M iterations), struggling to handle dynamic out-of-distribution scenarios, such as humans and animals, as illustrated in Figure 9. We should clarify that our baseline, Wan-I2V with only Plücker ray, suffers from inferior camera movements. While this setting achieves a strong VBench overall score, it compromises with poor camera metrics. Overall, the proposed PCDContoller achieves the optimal balance between video quality and camera precision. By integrating both Plücker rays and point clouds, it further enhances the performance in challenging scenes featuring substantial viewpoint changes, as validated in Table 1 and Figure 10.
5.3 Results of Unified Camera and Human Motion Control
Table 2: Quantitative results of unified camera and human motion controls. We re-render SMPL-X and Hamer under new camera trajectories for the comparison of RealisDance-DiT [73] without explicitly incorporating point clouds or cameras. $\dagger$ means masking out the foreground point clouds.
@@ 5.3 Results of Unified Camera and Human Motion Control
Camera Human Score Consist Consist Quality Quality Flicker Smooth
RealisDance-DiT [73] $\checkmark$ 85.21 93.03 95.34 57.89 68.71 97.44 98.82 0.549 0.195 0.547
@@
Uni3C (T2V)$\dagger$ $\checkmark$ $\checkmark$ 83.34 88.45 91.45 57.45 69.84 95.21 97.63 0.296 0.098 1.167
Uni3C (T2V) $\checkmark$ $\checkmark$ 83.16 88.67 91.38 56.79 69.42 95.14 97.57 0.262 0.083 0.606
Uni3C (I2V) $\checkmark$ $\checkmark$ 83.43 89.45 93.05 57.25 67.28 95.70 97.86 0.251 0.093 0.490
@@
We have developed a new benchmark of 50 videos, each featuring a person performing challenging motions. For guidance, SMPL-X is extracted using GVHMR [48]. Each video is assigned three different types of camera trajectories, resulting in a total of 150 test cases. To ensure that the person remains within the camera’s viewpoint, we employ a follow shooting technique for all test cases, adjusting the camera’s position based on the movement of the SMPL-X center. These noisy and subtle movements further increase the difficulty of camera control. We follow the same camera metrics as mentioned in Section 5.2. To facilitate generalization for motion transfer, RealisDance-DiT is not specifically designed to perfectly recover the first frame aligned with the reference image. Consequently, we remove the metrics that heavily depend on the reference view (I2V subject and background) for fairness.
@@
We show the quantitative results in Table 2, while qualitative results are displayed in Figure 11. To the best of our knowledge, there are currently no publicly available models that effectively address the challenge of unified camera and human motion controls. Therefore, we compare Uni3C against the baseline of RealisDance-DiT [73] and various ablation versions of our model. Notably, while RealisDance-DiT, which focuses solely on human control, achieves the best visual quality, it struggles to produce accurate camera trajectories, resulting in poor camera metrics. In contrast, Uni3C shows good VBench scores alongside impressive camera metrics. Note that the proposed PCDController can also be generalized to the T2V model with comparable quality, featuring the robust generalization of PCDController. Moreover, an interesting insight is revealed from Figure 11 that the aligned SMPL-X characters can further strengthen the camera controllability, while the PCDController trained with I2V formulation also enhances the visual quality and consistency of RealisDance-DiT. This illustrates the complementary features of these two components. Additionally, we show that the proposed Uni3C enables control of detailed hand motions under various camera trajectories as in Figure 12.
5.4 Ablation Study and Exploratory Discussions
Figure 6: Ablation results of Plücker ray and point clouds during the training phase. Point clouds enjoy highly accurate camera control against Plücker ray.

### v2

@@ Discussion.
We empirically find that our method demonstrates robust performance even when working with imperfect point clouds obtained through monocular depth unprojection. In this context, the point clouds serve as the primary camera control signal, facilitating the convergence of training rather than dominating the multi-view geometric and textural generations, as illustrated in Figure 4. Moreover, the lightweight PCDController retains its precise camera control capability over versatile fine-tuned Wan backbones, even without joint training. This flexibility allows for a range of downstream applications, showcasing the robustness of our approach.
4.2. Human Animation
In this paper, we explore the unified control through two human animation approaches, both of which are built on the Wan2.1 framework, targeting I2V and T2V, respectively. While these methods are not the primary focus of our work, we provide a brief introduction here. Formally, the concurrently pioneering work, RealisDance-DiT (Zhou et al., 2025b), directly replaces the Wan-I2V backbone for high-quality human animation during inference. RealisDance-DiT incorporates SMPL-X (Pavlakos et al., 2019) and Hamer (Pavlakos et al., 2024) as additional input conditions through newly zero-initialized layers to guide human motions, while camera-control features are added via the external PCDController as shown in Figure 2. To ensure the flexibility for motion transfer, RealisDance-DiT randomly selects the reference frame in the video sequence, which is not perfectly aligned with the given SMPL-X. RealisDance-DiT only trains self-attention modules and patchify encoders to confirm the generalization. However, we clarify that integrating the control branch trained within different backbones is still challenging, requiring a generalized control branch like our proposed PCDController. Furthermore, we tried another version, RealisDance-DiT-T2V, based on Wan-T2V without reference image conditions to explore the generalization of PCDController. Remarkably, PCDController adapts successfully to Wan-T2V, empowering it with impressive I2V ability.
4.3. Global 3D World Guidance
Figure 5. Overview of global 3D world guidance. (a) We first align SMPL-X characters from the human world space $W_{hum}$ to the environment world space $W_{env}$ with dense point clouds. (b) GeoCalib (Veicht et al., 2024) is used to calibrate the gravity direction of SMPL-X. (c) Rigid transformation coefficients $\tilde{s},\tilde{R},\tilde{t}$ are employed to align the whole SMPL-X sequence. We re-render all aligned conditions under specific camera trajectories as the global 3D world guidance.
@@ 4.3. Global 3D World Guidance
As illustrated in Figure 2, Uni3C adopts multi-modal conditions, including camera, point clouds, reference image, SMPL-X (Pavlakos et al., 2019), and Hamer (Pavlakos et al., 2024). The first three conditions are used for PCDController, while the latter three are used for human animations. We employ GVHMR (Shen et al., 2024) to recover SMPL-X characters. One can also retrieve desired motion sequences from motion datasets (Plappert et al., 2016; Punnakkal et al., 2021; Guo et al., 2022; Lin et al., 2023) or generate new motions through text-to-motion models (Jiang et al., 2023; Guo et al., 2024; Barquero et al., 2024). Although most of the conditions above are formulated as 3D presentations, they stay in two different world coordinates. We define the point cloud world coordinate as the environmental world space $W_{env}$, which is the main world space controlled by cameras, while the SMPL-X is placed in the human world space $W_{hum}$. It is non-trivial to control the camera across two different world spaces consistently. For example, determining the initial camera position within $W_{hum}$ is particularly ambiguous, especially for tasks involving motion transfer. Moreover, to facilitate flexible control over separate camera and human motions without conflicts, human conditions should be re-rendered under new camera trajectories. Therefore, we propose the global 3D world guidance that places the human condition into the “environment”, i.e., aligning SMPL-X from $W_{hum}$ to $W_{env}$ as shown in Figure 5(b).
@@
Fortunately, 2D human pose keypoints subtly bridge $W_{hum}$ and $W_{env}$. Formally, we first estimate 17 human keypoints $\{\mathbf{k}^{i}_{2D}\}_{i=1}^{17}\in\mathbb{R}^{17\times 2}$ from the reference view $I_{img}$ by ViTPose++ (Xu et al., 2023). Then, we unproject 2D keypoints into $W_{env}$ to obtain 3D keypoints $\{\mathbf{k}^{i}_{env}\}_{i=1}^{17}\in\mathbb{R}^{17\times 3}$ through metric monocular depth $\hat{D}_{img}$ and the intrinsic camera of the reference image. For SMPL-X in $W_{hum}$, the COCO17 regressor (Shen et al., 2024) is utilized to project the first frame’s SMPL-X character into $\{\mathbf{k}^{i}_{hum}\}_{i=1}^{17}$ corresponding to the same human keypoints. Consequently, a least-squares estimation (Umeyama, 1991) based rigid transformation can be used to align $\mathbf{k}_{hum}$ to $\mathbf{k}_{env}$ as follows:
(4) $\min_{\tilde{s},\tilde{R},\tilde{t}}\sum_{i=1}^{17}w_{i}\|(\tilde{s}\tilde{R}(\mathbf{k}^{i}_{hum})^{T}+\tilde{t})^{T}-\mathbf{k}^{i}_{env}\|^{2},$
@@
where $\tilde{s},\tilde{R},\tilde{t}$ indicate the optimized scaling factor, rotation matrix, and translation vector, respectively. $w_{i}$ denotes the confidence weight of 2D keypoint $\mathbf{k}^{i}_{2D}$. We discard any keypoints with confidence below 0.7, as they typically degrade alignment quality. Once the transformation parameters $\tilde{s},\tilde{R},\tilde{t}$ are determined, we apply them to all other SMPL-X sequences under the assumption that they share the same rigid transformation. However, even minor orientation errors can accumulate, leading to physically unrealistic motion trajectories, such as ascending into the sky or descending into the ground. To address this, we adopt GeoCalib (Veicht et al., 2024) to estimate the gravity direction in $W_{env}$, which is then employed to calibrate the SMPL-X to ensure parallel gravity directions, as illustrated in Figure 5(b). For the alignment of Hamer (Pavlakos et al., 2024), which shares common vertices with the hand parts of SMPL-X, Hamer can also be aligned to $W_{env}$ through the rigid transformation (Equation 4). Additionally, we address the issue of hand occlusion for Hamer by masking occluded hand parts based on the rendered depth from SMPL-X. After the alignments of SMPL-X and Hamer sequences, we place all conditions into $W_{env}$, establishing the global 3D world guidance that allows for rendering concurrently controlled conditions under arbitrary camera trajectories and human motions as shown in Figure 5(c). Finally, these re-rendered conditions are sent to PCDController and RealisDance-DiT for generated outcomes, as shown in Figure 2.
5. Experiments
5.1. Implementation Details
@@
PCDController is trained with the frozen Wan-I2V (Wang et al., 2025a) on multi-resolution images scaled from $480\times 768$, $512\times 720$, $608\times 608$, $720\times 512$, $768\times 480$] of 81 frames. The learning rate is warmed up to 1e-5 for 400 steps and then fixed. We train the model for 6,000 steps with batch size 32, while more training steps would slightly hurt Wan-I2V’s generalization. The training is accomplished with 64 H100 GPUs for 40 hours. We also provided results based on CogVideoX-5B-I2V (Yang et al., 2025), training for 20k steps with batch size 16. During the training, we randomly drop 10% texts, as well as 5% point cloud renderings and Plücker embeddings. For inference, we set the classifier-free guidance scale to 5.0 for textual conditions, keeping other guidance on the default scale 1.
@@
To ensure the generalization of PCDController, we collect large-scale training data for camera control, including DL3DV
 (Ling et al., 2024), RE10K (Zhou et al., 2018), ACID (Liu et al., 2021), Co3Dv2 (Reizenstein et al., 2021), Tartainair (Wang et al., 2020), Map-Free-Reloc (Arnold et al., 2022), WildRGBD (Xia et al., 2024), COP3D (Sinha et al., 2023), UCo3D (Liu et al., 2025). This comprehensive dataset encompasses various scenarios, featuring static and dynamic scenes, as well as object-level and scene-level environments. Furthermore, all datasets are annotated with metric-aligned monocular depth through the way proposed in (Cao et al., 2025) or are provided with ground-truth depth.
5.2. Results of Camera Control
Table 1. Quantitative results of camera control. VBench++ scores (%) are normalized (higher is better). Injected camera features are divided as Plücker ray and point clouds (Pcd). $\dagger$ denotes the results with challenging 360 camera rotations. Results of MotionCtrl and CameraCtrl are tested with the I2V version re-trained by (Zheng et al., 2025a). Methods in bold are the final setting of PCDController.
@@ 5.2. Results of Camera Control
MotionCtrl (Wang et al., 2024b) 82.48 89.42 91.64 54.99 55.02 91.93 95.85 89.92 91.11 0.345 0.263 2.547
CameraCtrl (He et al., 2025a) $\checkmark$ 83.69 90.17 92.21 55.71 54.13 93.32 96.65 93.34 94.01 0.354 0.239 2.306
CamI2V (Zheng et al., 2025a) $\checkmark$ 86.52 91.45 92.65 60.91 66.22 93.34 97.01 95.17 95.45 0.322 0.247 1.653
ViewCrafter (Yu et al., 2024) $\checkmark$ 85.39 89.69 91.68 55.13 64.33 92.94 97.66 95.59 96.11 0.210 0.117 0.873
SEVA (Zhou et al., 2025a) $\checkmark$ 87.39 91.86 93.36 56.79 68.43 95.74 98.59 97.08 97.23 0.077 0.029 0.223
@@
Ours (Wan-I2V) $\checkmark$ $\checkmark$ 88.27 92.20 93.37 58.99 71.96 95.56 98.66 97.38 98.01 0.102 0.031 0.246
[…]
@@
To evaluate camera control ability, we built an out-of-distribution benchmark with 32 images across various domains, including text-to-image generation [^1] https://github.com/black-forest-labs/flux, real-world (Knapitsch et al., 2017; Barron et al., 2022), object-centric, and human scenes as in Figure 17. Each image has four distinct camera trajectories, resulting in 128 test samples. Moreover, we collect a larger benchmark of 500 samples, comprising 50 scene images from megascenes (Tung et al., 2024) and 50 human scenes collected from the Internet. For each image, 5 random trajectories are assigned. We used VBench++ (Huang et al., 2024) to evaluate the video quality, while absolute translation error (ATE), relative translation error (RPE), and relative rotation error (RRE) are used to verify the camera precision. We utilize VGGT (Wang et al., 2025b) to produce extrinsic cameras for the generated images, which are then evaluated against the predefined cameras after trajectory alignment.
@@
We present the quantitative results in Table 1, comparing our model with ViewCrafter (Yu et al., 2024), SEVA (Zhou et al., 2025a), and the CogVideoX (Yang et al., 2025) version of our framework. The qualitative outcomes are shown in Figure 7. Our experiments demonstrate that point clouds significantly enhance the controllability of both the Wan2.1 and CogVideoX, as verified by the improvement of ATE, RPE, and RRE. Although SEVA achieves precise camera trajectories, it requires massive training with static multi-view data (0.8M iterations), struggling to handle dynamic out-of-distribution scenarios, such as humans and animals, as illustrated in Figure 7. We should clarify that our baseline, Wan-I2V with only Plücker ray, suffers from inferior camera movements. While this setting achieves a strong VBench overall score, it compromises with poor camera metrics. Overall, the proposed PCDController achieves the optimal balance between video quality and camera precision. By integrating both Plücker rays and point clouds, it further enhances the performance in challenging scenes featuring substantial viewpoint changes, as validated in Table 1 and Figure 8. Furthermore, our method also performs well in the large benchmark on megascene and Internet human images as listed in Table 2.
Table 2. Camera control results on large benchmark with 500 samples collected from megascenes (Tung et al., 2024) and Internet human scenes. ‘Overall’ means the average score across all metrics of VBench++.
@@
CamI2V (Zheng et al., 2025a) 83.62 0.8081 0.6858 2.5803
ViewCrafter (Yu et al., 2024) 84.26 0.6224 0.4098 1.2642
SEVA (Zhou et al., 2025a) 84.83 0.2473 0.0832 0.5329
PCDController 86.62 0.2613 0.0828 0.4612
5.3. Results of Unified Camera and Human Motion Control
Table 3. Quantitative results of unified camera and human motion controls. “Aligned” control means re-rendering human conditions under new camera trajectories in the environmental world space. $\dagger$ denotes masking out the foreground point clouds of humans. Method in bold is the final setting of Uni3C.
[…]
Camera Human Aligned Score Consist Consist Quality Quality Flicker Smooth
CamAnimate (Wang et al., 2024a) $\checkmark$ $\checkmark$ 82.45 89.20 90.52 57.42 67.40 93.94 96.20 0.619 0.419 2.035
RealisDance-DiT (Zhou et al., 2025b) $\checkmark$ $\checkmark$ 85.21 93.03 95.34 57.89 68.71 97.44 98.82 0.549 0.195 0.547
@@ 5.3. Results of Unified Camera and Human Motion Control
[…]
Uni3C (T2V) $\checkmark$ $\checkmark$ $\checkmark$ 83.16 88.67 91.38 56.79 69.42 95.14 97.57 0.262 0.083 0.606
Uni3C (I2V, unaligned) $\checkmark$ $\checkmark$ 82.98 89.16 92.40 56.02 66.78 95.65 97.84 0.270 0.111 0.639
Uni3C (I2V) $\checkmark$ $\checkmark$ $\checkmark$ 83.43 89.45 93.05 57.25 67.28 95.70 97.86 0.251 0.093 0.490
Figure 6. Results of unified camera and human motion controls. Leftmost images are reference views; the first row indicates aligned 3D world guidance.
@@
We have developed a new benchmark of 50 videos of 720p, each featuring a person performing challenging motions. For guidance, SMPL-X is extracted using GVHMR (Shen et al., 2024). Each video is assigned three different types of camera trajectories, resulting in a total of 150 test cases. To ensure that the person remains within the camera’s viewpoint, we employ a follow shooting technique for all test cases, adjusting the camera’s position based on the movement of the SMPL-X center. These noisy and subtle movements further increase the difficulty of camera control. We follow the same camera metrics as mentioned in Section 5.2. To facilitate generalization for motion transfer, RealisDance-DiT is not specifically designed to perfectly recover the first frame aligned with the reference image. Consequently, we remove the metrics that heavily depend on the reference view (I2V subject and background) in Table 3 for fairness. Note that our method performs well in 720p generation, even though it was trained using 480p data.
@@
We show quantitative results in Table 3, while qualitative results are displayed in Figure 6 and Figure 9. To our knowledge, there are currently very few available models that effectively address the challenge of unified camera and human motion controls. We begin by comparing to CamAnimate (Wang et al., 2024a), which enables simultaneous control of both conditions. However, due to the misalignment of body poses extracted relative to predefined camera trajectories, the results of CamAnimate suffer from conflicting human motions and backgrounds. Moreover, we evaluate Uni3C against the baseline of RealisDance-DiT (Zhou et al., 2025b) and various ablation versions of our model. Notably, while RealisDance-DiT, which focuses solely on human control, achieves the best visual quality, it struggles to produce accurate camera trajectories, resulting in poor camera metrics. In contrast, Uni3C shows good VBench scores alongside impressive camera metrics. Note that the proposed PCDController can also be generalized to the T2V model with comparable quality, featuring the robust generalization of PCDController. Moreover, an interesting insight is revealed from Figure 9 and Table 3 that the aligned SMPL-X characters can further strengthen the camera controllability, while the PCDController trained with I2V formulation also enhances the visual quality and consistency of RealisDance-DiT. This illustrates the complementary features of these two components. Additionally, we show that the proposed Uni3C enables control of detailed hand motions under various camera trajectories as in Figure 10.
5.4. Ablation Study and Exploratory Discussions